# Elemental Frequency-Based Supervised Classification Approach for the Search of Novel Topological Materials

### v1

@@ V CONCLUSION
In this study, we have used a frequency-based statistical approach for the creation of features to effectively implement machine learning techniques in the classification and discovery of non-trivial topological materials. The effectiveness of the suggested approach has been demonstrated by the classification accuracies of 82% and 83% for Support Vector Machine (SVM) and Random Forest (RF) models, respectively. We identified 22 common materials predicted by both models, which were further confirmed by density functional theory calculations. As an interpretable and computationally cheap substitute for traditional symmetry-based or structural-property-dependent descriptors, the frequency-based descriptors are derived solely from elemental occurrences. The models demonstrated dependable generalization, as confirmed by cross-validation and firstprinciplesfirst-principles computations, despite the small dataset. However, prediction accuracy is limited when material-specific physical or electronic properties are not included in the feature space. Model performance and generalizability could be significantly increased by expanding the feature space with more material descriptors, such as orbital characteristics, formation energy, or quantities derived from band structures. The findings confirm that a statistical frequency-based method offers a promising avenue for speeding up the identification of candidate topological materials when combined with machine learning and filtered using domain-specific criteria. To further increase the predictive ability of machine learning models for the discovery of quantum materials, future research may investigate hybrid feature spaces that combine statistical, structural, and electronic data.

### v2

@@ V CONCLUSION
In this study, we have used a frequency-based statistical approach for the creation of features to effectively implement machine learning techniques in the classification and discovery of non-trivial topological materials. The effectiveness of the suggested approach has been demonstrated by the classification accuracies of 82% and 83% for Support Vector Machine (SVM) and Random Forest (RF) models, respectively. We identified 22 common materials predicted by both models, which were further confirmed by density functional theory calculations. As an interpretable and computationally cheap substitute for traditional symmetry-based or structural-property-dependent descriptors, the frequency-based descriptors are derived solely from elemental occurrences. The models demonstrated dependable generalization, as confirmed by cross-validation and firstprinciplesfirst-principles computations, despite the small dataset. However, prediction accuracy is limited when material-specific physical or electronic properties are not included in the feature space. Model performance and generalizability could be significantly increased by expanding the feature space with more material descriptors, such as orbital characteristics, formation energy, or quantities derived from band structures. These findings confirmed that a statistical frequency-based method offers a promising avenue for speeding up the identification of candidate topological materials when combined with machine learning and filtered using domain-specific criteria. To further increase the predictive ability of machine learning models for the discovery of quantum materials, future research may investigate hybrid feature spaces that combine statistical, structural, and electronic data.